OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 2 points) Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

3(x + 1) = −2(x − 1) − 4 3x+3=-2x+2-4 3x+3=-2x-2 3x+2x=-2-3 5x=-5 x=-1